verify the identity
(cos(x)+isin(x))^3=cos(3x)+isin(3x)
so far i have completed the following
(cos(x)+isin(x))+(cos(x)+isin(x))+(cos(x)+isin(x))
(cos^2(x)+isin(x)cos(x)+isin(x)cos(x)+(-1)sin^2(x))(cos(x)+isin(x))
(cos^2(x)+2isin(x)cos(x)-sin^2(x))(cos(x)+isin(x))
cos^3(x)+isin(x)cos^2(x)+2isin(x)cos^2(x)+2(-1)sin^2(x)cos(x)-sin^2(x)cos(x)-isin^3(x)
cos^3(x)+3isin(x)cos^2(x)-3sin^2(x)cos(x)-isin^3(x)

Answer by Edwin McCravy(20056)


It's immediate by DeMoivre's theorem, but I suppose they want
you to prove it from scratch without using DeMoivre's theorem.  
OK, here's how I approach it:

[cos(x) + i·sin(x)]³ =

[cos(x) + i·sin(x)]²[cos(x) + i·sin(x)]

We'll work on the left bracketed expression only for awhile,
and just keep bringing down the right bracket expression
until we get the left bracket simplified:

[cos²(x) + 2i·cos(x)sin(x)+ i²·sin²(x)][cos(x) + i·sin(x)]

[cos²(x) + 2i·cos(x)sin(x)+ (-1)·sin²(x)][cos(x) + i·sin(x)]

[cos²(x) + 2i·cos(x)sin(x) - sin²(x)][cos(x) + i·sin(x)]

[{cos²(x) - sin²(x)} + i·{2cos(x)sin(x)}][cos(x) + i·sin(x)]

[{cos²(x) - sin²(x)} + i·{2sin(x)cos(x)}][cos(x) + i·sin(x)]

Use the identities:

cos²(A) - sin²(A) = cos(2A)
2sin(A)cos(A) = sin(2A)

[cos(2x) + i·sin(2x)][cos(x) + i·sin(x)] =

Now we finally use the right bracket and multiply out:

cos(2x)cos(x) + i·cos(2x)sin(x) + i·sin(2x)cos(x) + i²·sin(2x)cos(x) =

cos(2x)cos(x) + i·[cos(2x)sin(x) + sin(2x)cos(x)] + (-1)·sin(2x)cos(x) =

cos(2x)cos(x) + i·[cos(2x)sin(x) + sin(2x)cos(x)] - sin(2x)cos(x) =

cos(2x)cos(x) - sin(2x)sin(x) + i·[cos(2x)sin(x) + sin(2x)cos(x)] =

cos(2x)cos(x) - sin(2x)sin(x) + i·[sin(2x)cos(x) + cos(2x)sin(x)] =

Use the identities: 

cos(A)cos(B) - sin(A)sin(B) = cos(A+B)
sin(A)cos(B) + cos(A)sin(B) = sin(A+B)

cos(2x+x) + i·sin(2x+x)

cos(3x) + i·sin(3x)

You can probably leave out some of those steps where I just regrouped
terms.

Edwin